# Structural Asymmetry and Rigidity of Bounded Solutions for a Parametrized Complex Linear Differential Equation

### v4

@@ Abstract
In this manuscript, we introduce a family of parametrized nonhomogeneous linear complex differential equations on $1,\infty)$, depending on a complex parameter lying in the critical strip. We identify a suitable rigidity hypothesis under which a phenomenon of structural asymmetry is established: if two solutions with the same initial condition equal to 11, corresponding respectively to the parameters $s$ and $1-\bar{s}$, are both bounded on $1,+\infty)$, then $\Re(s)=\frac{1}{2}$.
@@ 1 Main results
In this manuscript, we study the following non-homogeneous linear complex parametrized differential equations
@@
$t\in[1,+\infty),\quad x(1)=z,\quad x:[1,+\infty)\to\mathbb{C},$
where $w\in\mathbb{C}$ and $\eta:[1,+\infty)\to\mathbb{C}$ is a locally integrable function and satisfy $|\eta(t)|<1$ for all $t\geq 1$. The equation (1) can be viewed as a differential equation defined on $1,+\infty)$ with values in $\mathbb{R}^{2}$, and we are interested in the initial conditions of the bounded solutions. This result provides an answer to the Dynamical Conjecture formulated in [2]. In the study that follows, we fix a specific function $\eta$ which is locally integrable function and satisfy $|\eta(t)|<1$ for all $t\geq 1$. The function
$t\mapsto t^{\Re(w)}\int_{1}^{t}u^{-1-w}\eta(u)\,du,\quad t\geq 1,$
is continuous on $1,+\infty)$. The differential equation (1) is a non-homogeneous linear differential equation. Then, for every $z\in\mathbb{C}$, there exists a unique continuous solution $\psi_{w}(z,\cdot)$ of (1) such that $\psi_{w}(z,1)=z$, which is given by
@@
Denote by $B\subset\mathbb{C}$ the following set
$B:=\Big\{w\in\mathbb{C}:\quad\Re(w)>0\Big\}.$
Let $\mu$ denote the function $\mu:B\mapsto\mathbb{C}$, defined as
$\mu(s)=-(1-s)\int_{1}^{+\infty}u^{-1-s}\eta(u)\,du,\quad\forall s\in B.$
In the following lemma, we are interested in bounded solutions on $1,+\infty)$ and their initial condition.
@@
Let $w\in B$ and $z\in\mathbb{C}$. Consider the continuous solution $\psi_{w}(z,\cdot)$ of the differential equation (1) as given in Notation 1. Then
@@
Let $w\in B$ be fixed. There is a unique bounded solution; all other solutions diverge in norm. Let $z\in\mathbb{C}$ and suppose that $\sup_{t\geq 1}|\psi_{w}(z,t)|<+\infty$. By equation (2), this implies
@@
which gives $z=\mu(w)$. To prove that $\sup_{t\geq 1}|\psi_{w}(\mu(w),t)|<+\infty$, observe that by definition of $\mu(w)$,
@@
$\sup_{t\geq 1}|\psi_{w}(\mu(w),t)|\leq\frac{|1-w|}{\Re(w)}\sup_{t\geq 1}|\eta(t)|<\frac{|1-w|}{\Re(w)}.$
From the perspective of the qualitative theory of dynamical systems, the differential equation (1) can be viewed as a differential equation with the parameter $\Re(w)$. Intuitively, this introduces a order structure within the set of solutions with a same initial condition and parameterized by $\Re(w)$. We show this in the following proposition.
@@
Let be $s:\sigma+i\tau\in B$. Consider the continuous solutions $\psi_{s}(1,\cdot)$ and $\psi_{1-\bar{s}}(1,\cdot)$ of the differential equation (1) as given in Notation 1. The function $\delta_{s}$ defined as
@@
satisfies the following two equalities:
$\dot{\delta}_{s}(1)$ $=(2\sigma-1)\Big(1-\Re\big(\eta(1)\big)\Big),$
${\delta}_{s}(t)$ $=\frac{1}{2}\int_{1}^{t}u^{-1}\delta_{s}(u)du+(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-\frac{1}{2}}\int_{1}^{u}v^{-\frac{3}{2}}{\delta}_{s}(v)\,dv\,du$
$\ +4(\sigma-\frac{1}{2})\Re\!\left(\psi_{\frac{1}{2}+i\,\tau}(1,t)-\psi_{i\tau}(1,t)\right),\qquad\forall t\geq 1.$
According to the notation 1 consider the following rigidity hypothesis H:
  • H.
    For every $\beta\in\mathbb{R}$ and every $t>1$,
    $\Re\!\left(\psi_{\frac{1}{2}+i\beta}(1,t)-\psi_{i\beta}(1,t)\right)>0.$
The previous hypothesis suggest the rigidity of the sign. It is suitable in order to apply the results of this dynamical study to the Riemann zeta function.
The following theorem shows that, under the rigidity hypothesis H, if $\Re(s)\neq\tfrac{1}{2}$, then the continuous solutions $\psi_{s}(1,\cdot)$ and $\psi_{1-\bar{s}}(1,\cdot)$ of the differential equation (1) cannot both be bounded on $1,+\infty)$.
@@
Consider the function $\mu$ defined in Notation 2. Under hypothesis H, for every $s\in\mathbb{C}$ with $\Re(s)\in(0,1)/\{\frac{1}{2}\}$, we have
@@ 2 Proof of the Proposition 4 and Theorem 5
Let be $w\in B$ and denote $w:=\alpha+i\beta$. For every $t\geq 1$ denote
@@
Let be $s:=\sigma+i\tau\in B$ such that $\sigma\in(\frac{1}{2},1)$. From the previous equation, we have
@@
Define
By the notation (4), we have ${x_{s}}(t)-{x}_{1-\bar{s}}(t):=t^{-\frac{1}{2}}\delta_{s}(t)$. We obtained
$\dot{\delta}_{s}(t)=$ $\frac{1}{2}t^{-1}\delta_{s}(t)+(\sigma-\frac{1}{2})^{2}t^{-\frac{1}{2}}\int_{1}^{t}u^{-\frac{3}{2}}\delta_{s}(u)\,du$
$+2(\sigma-\frac{1}{2})t^{-\frac{1}{2}}\Big(1-t^{-\frac{1}{2}}\Re\big(\,t^{-i\tau}\,\eta(t)\big)\Big)$
$+2(\sigma-\frac{1}{2})t^{-\frac{1}{2}}\Re\Big((\frac{1}{2}-i\tau)\int_{1}^{t}u^{-\frac{3}{2}-i\tau}\,\eta(u)\,du\Big).$
Since ${\delta}_{s}(1)=0$ and since by hypothesis we have $|\eta(1)|<1$, then
$\dot{\delta}_{s}(1)=(2\sigma-1)\Big(1-\Re\big(\eta(1)\big)\Big).$
Now, prove the second item of the present Proposition. Integrate the previous equation, we obtain
${\delta}_{s}(t)=\frac{1}{2}\int_{1}^{t}u^{-1}\delta_{s}(u)du+(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-\frac{1}{2}}\int_{1}^{u}v^{-\frac{3}{2}}\dot{\delta}_{s}(v)\,dv\,du$
$+2(\sigma-\frac{1}{2})\int_{1}^{t}u^{-\frac{1}{2}}\Re\Big(1-u^{-\frac{1}{2}-i\tau}\,\eta(u)\Big)du$
$+2(\sigma-\frac{1}{2})\Re\Big((\frac{1}{2}-i\tau)\int_{1}^{t}u^{-\frac{1}{2}}\int_{1}^{u}v^{-\frac{3}{2}-i\tau}\,\eta(v)\,dv\,du\Big).$
Use the integration by pars formula to the last integral, we get
${\delta}_{s}(t)=\frac{1}{2}\int_{1}^{t}u^{-1}\delta_{s}(u)du+(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-\frac{1}{2}}\int_{1}^{u}v^{-\frac{3}{2}}\dot{\delta}_{s}(v)\,dv\,du$
$+2(\sigma-\frac{1}{2})\int_{1}^{t}u^{-\frac{1}{2}}\Re\Big(1-u^{-\frac{1}{2}-i\tau}\,\eta(u)\Big)du$
$+4(\sigma-\frac{1}{2})t^{\frac{1}{2}}\Re\Big((\frac{1}{2}-i\tau)\int_{1}^{t}v^{-\frac{3}{2}-i\tau}\,\eta(v)\,dv\Big)$
$-4(\sigma-\frac{1}{2})\Re\Big((\frac{1}{2}-i\tau)\int_{1}^{t}u^{-1-i\tau}\,\eta(u)\,du\Big).$
Equivalent to
${\delta}_{s}(t)=$ $\frac{1}{2}\int_{1}^{t}u^{-1}\delta_{s}(u)du+(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-\frac{1}{2}}\int_{1}^{u}v^{-\frac{3}{2}}{\delta}_{s}(v)\,dv\,du$
$+4(\sigma-\frac{1}{2})t^{\frac{1}{2}}\big[(\frac{1}{2}-i\tau)\int_{1}^{t}v^{-\frac{3}{2}-i\tau}\,\eta(v)\,dv\Big]$
$-4(\sigma-\frac{1}{2})\Big[1+\Re\Big((1-i\tau)\int_{1}^{t}u^{-1-i\tau}\,\eta(u)\,du\Big)\Big].$
According to Notation 1, we obtain
${\delta}_{s}(t)=$ $\frac{1}{2}\int_{1}^{t}u^{-1}\delta_{s}(u)du+(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-\frac{1}{2}}\int_{1}^{u}v^{-\frac{3}{2}}{\delta}_{s}(v)\,dv\,du$
$+4(\sigma-\frac{1}{2})\Re\!\left(\psi_{\frac{1}{2}+i\beta}(1,t)-\psi_{i\beta}(1,t)\right).$
@@
Let be $s:=\sigma+i\tau\in B$ such that $\sigma\in(0,1)/\{\frac{1}{2}\}$. By symmetry, we may assume that $\sigma\in(\frac{1}{2},1)$. For every $t\geq 1$ denote
@@
Thank to the Proposition 4, we have $\dot{\delta}_{s}(1)>0$. By continuity there exists $\epsilon>0$ such that
@@
${\delta}_{s}(t)=$ $\frac{1}{2}\int_{1}^{t}u^{-1}\delta_{s}(u)du+(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-\frac{1}{2}}\int_{1}^{u}v^{-\frac{3}{2}}{\delta}_{s}(v)\,dv\,du$
$+4(\sigma-\frac{1}{2})\Re\!\left(\psi_{\frac{1}{2}+i\,\tau}(1,t)-\psi_{i\tau}(1,t)\right).$
Since $2\sigma-1>0$, thanks to the hypothesis H, we obtain
${\delta}_{s}(t)>\frac{1}{2}\int_{1}^{t}u^{-1}\delta_{s}(u)du+(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-\frac{1}{2}}\int_{1}^{u}v^{-\frac{3}{2}}{\delta}_{s}(v)\,dv\,du,\quad\forall t>1.$ (8)
@@
${\delta}_{s}(t_{*})>\frac{1}{2}\int_{1}^{t_{*}}u^{-1}\delta_{s}(u)du+(\sigma-\frac{1}{2})^{2}\int_{1}^{t_{*}}u^{-\frac{1}{2}}\int_{1}^{u}v^{-\frac{3}{2}}{\delta}_{s}(v)\,dv\,du,$
@@
$\dot{\delta}_{s}(t)>(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-\frac{1}{2}}\int_{1}^{2}v^{-\frac{3}{2}}{\delta}_{s}(v)\,dv\,du,\quad\forall t>1.$
That gives
${\delta}_{s}(t)>2(\sigma-\frac{1}{2})^{2}(t^{\frac{1}{2}}-1)\int_{1}^{2}u^{-\frac{3}{2}}{\delta}_{s}(u)du,\quad\forall t>1.$
@@
More precisely, the proof of Proposition 4 shows that under the rigidity hypothesis H, for all $s\in B$ with $\Re(s)\in(\frac{1}{2},1)$, the function $\mu$ satisfies the following equation
$\Re\big(\mu(s)\big)-\Re\big(\mu(1-\overline{s})\big)\leq 2\Re\big(\eta(1)\big).$
In this case, $\dot{\delta}_{s}(1)\leq 0$; That is, the function does not necessary start positively from 11. Hence the proof of the Theorem 5 cannot be continued.

### v6

# Rigidity Hypothesis for the Structural Asymmetry of the Bounded Solutions of a Parametric Complex Linear Differential Equation

@@ Abstract
In this manuscript, we introduce a family of parametrized non-homogeneous linear complex differential equations on $1,\infty)$, depending on a complex parameter lying in the critical strip. We identify a rigidity hypothesis along the critical line, which establishes a structural asymmetry: if two solutions with the same initial condition equal to 11, corresponding respectively to the parameters $s$ and $1-\bar{s}$, are both bounded on $1,+\infty)$, then $\Re(s)=\frac{1}{2}$. This rigidity hypothesis is satisfied when the non-homogeneous term is the fractional part function.
@@ 1 Main results
In this manuscript, we study the following non-homogeneous linear complex differential equations:
$\dot{\theta}=wt^{-1}\theta+t^{-1}\eta(t),\quad\theta:[1,+\infty)\to\mathbb{C},$
where $w\in\mathbb{C}$ is the parameter and the non-homogeneous term $\eta$ is a normalized, locally integrable function satisfying $|\eta(t)|<1$ for all $t\geq 1$. We are interested in the initial conditions of the bounded solutions. To this end, we analyze the behavior of the transformed solution $t\mapsto(1-w)t^{-i\Im(w)}\theta(t)$, which leads us to consider the following differential equation:
@@
$t\in[1,+\infty),\quad x(1)=z,\quad x:[1,+\infty)\to\mathbb{C}.$
In the following study, we fix a specific function $\eta$ as defined above. The function
$t\mapsto t^{\Re(w)}\int_{1}^{t}u^{-1-w}\eta(u)\,du,\quad t\geq 1.$
is absolutely continuous on $1,+\infty)$. The differential equation (1) is a non-homogeneous linear differential equation. Then, for every $z\in\mathbb{C}$, there exists a unique continuous solution $\psi_{w}(z,\cdot)$ of (1) such that $\psi_{w}(z,1)=z$, which is given by
@@
Denote by $\mathbb{C}_{+}$ the right half-plane defined as
$\mathbb{C}_{+}:=\Big\{w\in\mathbb{C}:\quad\Re(w)>0\Big\}.$
Let $\mu$ denote the function $\mu:\mathbb{C}_{+}\mapsto\mathbb{C}$, defined as
$\mu(s)=-(1-s)\int_{1}^{+\infty}u^{-1-s}\eta(u)\,du,\quad\forall s\in\mathbb{C}_{+}.$
The function $\mu$ is defined for all $s\in\mathbb{C}_{+}$. Indeed, the integral is absolutely convergent, since for all $w\in\mathbb{C}_{+}$ we have $\Re(w)>0$. In the following lemma, we are interested in bounded solutions on $1,+\infty)$ and their initial condition.
@@
Let $w\in\mathbb{C}_{+}$ and $z\in\mathbb{C}$. Consider the continuous solution $\psi_{w}(z,\cdot)$ of the differential equation (1) as given in Notation 1. Then
@@
Let $w\in\mathbb{C}_{+}$ be fixed. There is a unique bounded solution; all other solutions diverge in norm. Let $z\in\mathbb{C}$ and suppose that $\sup_{t\geq 1}|\psi_{w}(z,t)|<+\infty$. By equation (2), this implies
@@
which gives $z=\mu(w)$. Now, prove that that $\sup_{t\geq 1}|\psi_{w}(\mu(w),t)|<+\infty$. By definition of $\mu(w)$, we have
@@
$\sup_{t\geq 1}|\psi_{w}(\mu(w),t)|\leq\frac{|1-w|}{\Re(w)}\sup_{t\geq 1}|\eta(t)|\leq\frac{|1-w|}{\Re(w)}.$
The differential equation (1) can be viewed as a differential equation with the parameter $\Re(w)$. Intuitively, this introduces a order structure within the set of solutions with a same initial condition and parameterized by $\Re(w)$. We start to exhibit this order between solutions through the following proposition.
@@
Consider the continuous solutions of the differential equation (1) as given in Notation 1. For every $s:=\sigma+i\,\tau\in\mathbb{C}_{+}$ with $\sigma\in(0,1)$, the function $\delta_{s}$ defined as
@@
satisfies the following integral equation:
$\delta_{s}(t)=$ $(\sigma-\frac{1}{2})^{2}t^{\frac{1}{2}}\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-\frac{3}{2}}\delta_{s}(v)\,dv\,du$
$+(2\sigma-1)t^{\frac{1}{2}}\int_{1}^{t}u^{-\frac{3}{2}}\Re\big(\psi_{\frac{1}{2}+i\,\tau}(1,u)\big)\,du$
$+(2\sigma-1)t^{\frac{1}{2}}\Re\Big(\frac{1-t^{-\frac{1}{2}}\psi_{\frac{1}{2}+i\,\tau}(1,t)}{\frac{1}{2}-i\,\tau}\Big),\qquad\qquad\forall t\geq 1.$
Denote by $L$ the critical ligne difened as
$L:=\{w\in\mathbb{C},\quad\Re(w)=\frac{1}{2}\}.$
Given the previous proposition, we consider the following rigidity hypothesis:
$\forall w\in L,\ \forall t\geq 1:\quad\Re\Big(\frac{1-t^{-\frac{1}{2}}\psi_{w}(1,t)}{1-w}+\int_{1}^{t}u^{-\frac{3}{2}}\psi_{w}(1,u)\,du\Big)\geq 0.$ (H)
The previous hypothesis represents a rigidity of the sign along the critical line. It is suitable in order to apply the results of this dynamical study to the Riemann zeta function.
The following theorem shows that, under the rigidity hypothesis (H), if $\Re(s)\neq\tfrac{1}{2}$, then the continuous solutions $\psi_{s}(1,\cdot)$ and $\psi_{1-\bar{s}}(1,\cdot)$ of the differential equation (1) cannot both be bounded on $1,+\infty)$.
@@
Consider the function $\mu$ given by Notation 2. Under hypothesis (H), for every $s\in\mathbb{C}$ with $\Re(s)\in(0,1)/\{\frac{1}{2}\}$, we have
@@ 2 Proof of the Proposition 4 and Theorem 5
Let be $w\in\mathbb{C}_{+}$ and denote $w:=\alpha+i\,\beta$. For every $t\geq 1$ denote
@@
Let be $s:=\sigma+i\,\tau\in\mathbb{C}_{+}$ such that $\sigma\in(0,1)$. From the previous equation, we have
@@
Use the fact that ${x}_{s}(1)-{x}_{1-\bar{s}}(1)=0$ and integrate, we obtain
${x}_{s}(t)-{x}_{1-\bar{s}}(t)=$ $(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-1}\big({x_{s}}(v)-{x}_{1-\bar{s}}(v)\big)\,dv\,du$
$+2(\sigma-\frac{1}{2})\Re\big(\,\int_{1}^{t}u^{-1}\Big(1-u^{-\frac{1}{2}}u^{-i\,\tau}\,\eta(u)\big)du\Big)$
$+2(\sigma-\frac{1}{2})\Re\Big((\frac{1}{2}-i\tau)\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-\frac{3}{2}-i\,\tau}\,\eta(v)\,dv\,du\Big).$
From the equation (2), for every $\beta\in\mathbb{R}$ and $t\geq 1$, we have
$\psi_{\frac{1}{2}+i\,\beta}(1,t)=t^{\frac{1}{2}}\Big[1+(\frac{1}{2}-i\beta)\int_{1}^{t}u^{-\frac{3}{2}-i\,\beta}\,\eta(u)\,du\Big],\quad\forall\beta\in\mathbb{R}.$
Therefore,
${x}_{s}(t)-{x}_{1-\bar{s}}(t)=$ $(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-1}\big({x_{s}}(v)-{x}_{1-\bar{s}}(v)\big)\,dv\,du$
$+2(\sigma-\frac{1}{2})\int_{1}^{t}u^{-\frac{3}{2}}\Re\Big(\psi_{\frac{1}{2}+i\,\tau}(1,u)-u^{-i\,\tau}\,\eta(t)\Big)\,du,$
and that can be written as
${x}_{s}(t)-{x}_{1-\bar{s}}(t)$ $=(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-1}\big({x_{s}}(v)-{x}_{1-\bar{s}}(v)\big)\,dv\,du$
$+(2\sigma-1)\int_{1}^{t}u^{-\frac{3}{2}}\Re\big(\psi_{\frac{1}{2}+i\,\tau}(1,u)\big)\,du$
$+(2\sigma-1)\Re\Big(\frac{1-t^{-\frac{1}{2}}\psi_{\frac{1}{2}+i\,\tau}(1,t)}{\frac{1}{2}-i\,\tau}\Big).$
By the notation (4), we have
${x_{s}}(t)-{x}_{1-\bar{s}}(t)=t^{-\frac{1}{2}}\delta_{s}(t):=t^{-\frac{1}{2}}\Re\big({\psi_{s}}(1,t)-{\psi}_{1-\bar{s}}(1,t)\big).$
@@
Let be $s:=\sigma+i\,\tau\in\mathbb{C}$ such that $\sigma\in(0,1)/\{\frac{1}{2}\}$. By symmetry, we may assume that $\sigma\in(\frac{1}{2},1)$. For every $t\geq 1$ denote
@@
From the equation (1) , we have
$\dot{\delta}_{s}(t)=t^{-1}\Re\Big(\sigma\psi_{s}(1,t)-(1-\sigma)\psi_{1-\bar{s}}(1,t)\Big)-(2\sigma-1)\,t^{-1}\Re\big(t^{-i\,\tau}\eta(t)\Big),$
Since $\psi_{s}(1,1)=\psi_{1-\bar{s}}(1,1)=1$ then ${\delta}_{s}(1)=0$. By hypothesis, we have $|\eta(t)|<1$ for all $t\geq 1$. Since $\sigma\in(\frac{1}{2},1)$ the previous differential equation implies $\dot{\delta}_{s}(1)>0$. By continuity there exists $\epsilon>0$ such that
@@
$\delta_{s}(t)=$ $(\sigma-\frac{1}{2})^{2}t^{\frac{1}{2}}\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-\frac{3}{2}}\delta_{s}(v)\,dv\,du$
$+(2\sigma-1)t^{\frac{1}{2}}\int_{1}^{t}u^{-\frac{3}{2}}\Re\big(\psi_{\frac{1}{2}+i\,\tau}(1,u)\big)\,du$
$+(2\sigma-1)t^{\frac{1}{2}}\Re\Big(\frac{1-t^{-\frac{1}{2}}\psi_{\frac{1}{2}+i\,\tau}(1,t)}{\frac{1}{2}-i\,\tau}\Big),\qquad\qquad\forall t\geq 1.$
Since $2\sigma-1>0$, thanks to the hypothesis (H), we obtain
${\delta}_{s}(t)\geq(\sigma-\frac{1}{2})^{2}t^{\frac{1}{2}}\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-\frac{3}{2}}\delta_{s}(v)\,dv\,du,\quad\forall t\geq 1.$ (8)
@@
${\delta}_{s}(t_{*})\geq(\sigma-\frac{1}{2})^{2}t_{*}^{\frac{1}{2}}\int_{1}^{t_{*}}u^{-1}\int_{1}^{u}v^{-\frac{3}{2}}\delta_{s}(v)\,dv\,du.$
@@
${\delta}_{s}(t)>(\sigma-\frac{1}{2})^{2}t^{\frac{1}{2}}\int_{1}^{2}u^{-1}\int_{1}^{u}v^{-\frac{3}{2}}\delta_{s}(v)\,dv\,du,\quad\forall t>1.$
@@
Remark 6.
  • Under hypothesis (H), the process used in the proofs of Proposition 4 and Theorem 5 shows that: For all $s\in\mathbb{C}$ with $\Re(s)\in(\frac{1}{2},1)$, the function $\mu$ satisfies the following implication
    $\Re\Big(\mu(s)+\mu(1-\overline{s})\Big)\geq 2\implies\Re\Big(\mu(s)-\mu(1-\overline{s})\Big)<0.$
    In this case, the function ${\delta}_{s}$ start with a negative values at 11. Hence, the proof of Theorem 5 cannot be continued.
  • We recall that the rigidity hypothesis (H) is given as:
    $\forall w\in L,\ \forall t\geq 1:\quad\Re\Big(\frac{1-t^{-\frac{1}{2}}\psi_{w}(1,t)}{1-w}+\int_{1}^{t}u^{-\frac{3}{2}}\psi_{w}(1,u)\,du\Big)\geq 0.$
    From the equation (2), for every $\beta\in\mathbb{R}$ and $t\geq 1$, we have
    $\psi_{\frac{1}{2}+i\,\beta}(1,t)=t^{\frac{1}{2}}\Big[1+(\frac{1}{2}-i\beta)\int_{1}^{t}u^{-\frac{3}{2}-i\,\beta}\,\eta(u)\,du\Big],$
    Hence, the rigidity hypothesis (H) is equivalent to:
    $\forall\beta\in\mathbb{R},\quad\forall t\geq 1:\quad\Re\big(f_{\beta}(t)\big)\geq 0,$
    where
    $f_{\beta}(t):=$ $\ln(t)-\int_{1}^{t}u^{-\frac{3}{2}-i\,\beta}\,\eta(u)\,du$
    $+(\frac{1}{2}-i\beta)\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-\frac{3}{2}-i\,\beta}\,\eta(v)\,dv\,du.$